OpenStudy (anonymous):

product of a number d and 16

OpenStudy (anonymous):

someone helppppppppppp

OpenStudy (anonymous):

product means what?

OpenStudy (anonymous):

didn't say the question is find the variable expression for word phrase the product of a number d and 16

OpenStudy (anonymous):

i know that, do you know what that means? Again, do you know what product means?

OpenStudy (anonymous):

noo

OpenStudy (anonymous):

really? product means you multiply two numbers.

OpenStudy (anonymous):

so the answer d times 16?????

OpenStudy (anonymous):

i think

OpenStudy (anonymous):

never mind ill give up

jimthompson5910 (jim_thompson5910):

`so the answer d times 16?????` yep that's correct @romans116

jimthompson5910 (jim_thompson5910):

you can say any of the following `d times 16` `d * 16` `16*d` `16d` all say the same thing. The star or asterisk `*` symbol is the keyboard symbol that means multiply. When you have a number (16) and a letter (d) put together like you see in `16d`, there is an implicit multiplication symbol between the letter and number

OpenStudy (anonymous):

ok so h increased by 10 would be h+10

pooja195 (pooja195):

Yep :)

jimthompson5910 (jim_thompson5910):

`ok so h increased by 10 would be h+10` agreed, you have the right answer

pooja195 (pooja195):

increased = to raise or add

pooja195 (pooja195):

so ye thats right :)

OpenStudy (anonymous):

sum of a and b is a+b

OpenStudy (anonymous):

and the product of m and n=m+n two less than k is=2-k s divided by n is s /n

jimthompson5910 (jim_thompson5910):

`sum of a and b is a+b` yes

jimthompson5910 (jim_thompson5910):

`product of m and n=m+n` incorrect